OpenStudy (anonymous):

I do award metals to those who help! I just need someone to tell me if my answers are right please!

OpenStudy (anonymous):

1) Which of the following atoms contains the same number of neutrons as an atom of Fluorine-19? neon-20 oxygen-19<--- sodium-22 nitrogen-16

OpenStudy (anonymous):

@nechirwan do you know if that one is correct?

OpenStudy (anonymous):

@helpout3 @HelpBlahBlahBlah do you know if this one is correct?

OpenStudy (nechirwan):

Fluorine: 19-9=10 Neon: 20-10=10 Oxygen: 19-8=11 Sodium: 22-11=11 Nitrogen: 16-7= 9 So Fluorine-19 has the same amount of neutrons as Neon-20

OpenStudy (anonymous):

I have a question, how do you know what to subtract? like how do you know to subtract 9 from 19, 10 from 20, etc?

OpenStudy (nechirwan):

19 is number of proton +neutron) - the number of proton u u get the number of neutron

OpenStudy (anonymous):

oh okay, got it. Thank you :-) I have another one

OpenStudy (anonymous):

Which of the following contains the greatest number of moles? 9.1x10^23 atoms Ag 4.7 mol Ag 1.3 mol fe 56 g fe

OpenStudy (anonymous):

I am thinking A

OpenStudy (nechirwan):

what is mole rule?

OpenStudy (nechirwan):

sorry the rule is wrong

OpenStudy (anonymous):

it's okay. But I forget the mole rule

OpenStudy (nechirwan):

mole = mo. of atom div by Avogadro number

OpenStudy (anonymous):

that looks so confusing lol